Need help on probability.
1) A doughnut factory made 113 doughnuts, jelly was filled in 17 doughnuts. Based on this sample of values, how many doughnuts will be filled expected out of 1700 doughnuts?

Answer by stanbon(75887)

A doughnut factory made 113 doughnuts, jelly was filled in 17 doughnuts. Based on this sample of values, how many doughnuts will be filled expected out of 1700 doughnuts?
-----------
x/1700 = 17/113
x= (1700)(17/113
x=255.75
Rounding down: x=255
====================
Cheers,
Stan H.